OpenStudy (anonymous):

a-b-c A=4 B=-8 and C=-5

OpenStudy (anonymous):

so it would be 4- -8 - -5 the negative switches with subtration to form positive making it 4+8+5 so your answer would be 17

OpenStudy (anonymous):

4-(-8)=12-(-5)=17

OpenStudy (anonymous):

okay thanks very much(:

OpenStudy (anonymous):

the answer is 7